### v3

# Mechanical hysterons with tunable interactions of general sign

Abstract
Hysterons are elementary units of hysteresis that underlie many complex behaviors of non-equilibrium matter. Because models of interacting hysterons can describe disordered matter, this suggests that artificial systems could respond to mechanical inputs in precise and targeted ways. Specifying the properties of hysterons and their interactions could thus be a general method for realizing arbitrary non-equilibrium behaviors. Elastic structures including slender beams, creased sheets, and shells are clear candidates for artificial hysterons, but complete control of their interactions has seemed impractical or impossible. Here we report a mechanical hysteron composed of rigid bars and linear springs, which has controllable properties and tunable interactions of general sign that can be reciprocal or non-reciprocal. We derive a mapping from the system parameters to the hysteron properties, and we show how collective behaviors of multiple hysterons can be targeted by adjusting geometric parameters on the fly. By transforming an abstract hysteron model into a physical design platform, our work demonstrates a route toward designed materials that can sense, compute, and respond to their mechanical environment.
When an amorphous solid, a disordered magnet, or a crumpled sheet is subjected to an oscillating global drive, it navigates a complex pathway through a multitude of locally-stable states Paulsen and Keim (2024). A powerful approach for understanding these systems has been to focus on the localized hysteretic rearrangements—hysterons—that arise within them Preisach (1935); Falk and Langer (2011); Keim and Arratia (2014); Mungan et al. (2019); Mungan and Terzi (2019); Terzi and Mungan (2020); Keim et al. (2020); Keim and Medina (2022); Shohat et al. (2023); Muhaxheri and Santangelo (2024); Martínez-Calvo et al. (2024). In these abstract hysteron models, each hysteron $i$ possesses a pair of thresholds $\gamma_{i}^{+}$ and $\gamma_{i}^{-}$ at which it switches between its two possible states, $s_{i}=\pm 1$. More general models incorporate interactions between hysterons Keim and Paulsen (2021); Lindeman and Nagel (2021); van Hecke (2021); Szulc et al. (2022), so that the switching thresholds $\gamma_{i}^{\pm}$ depart from their “bare” values $\gamma_{i,0}^{\pm}$ as follows:
$\gamma_{i}^{\pm}(S)=\gamma_{i,0}^{\pm}-\sum_{j\neq i}J_{ij}s_{j}\ ,$ (1)
where $S$ is a microstate that specifies the $\{s_{j}\}_{j\neq i}$ and the $J_{ij}$ encode cooperative ($J_{ij}>0$) or frustrated ($J_{ij}<0$) interactions that may be non-reciprocal ($J_{ij}\neq J_{ji}$).
In addition to offering a foothold for understanding disordered media, Eq. 1 suggests a design principle for a new class of synthetic matter. If the thresholds and interactions can be designed, materials could be created with tailored responses to different mechanical environments. As the states $s_{i}$ are readily interpreted as bits, such materials are endowed with digital memory and the ability to perform computations Yasuda et al. (2021). Several mechanical metamaterials have been developed to this end, including origami bellows Jules et al. (2022), corrugated sheets Bense and van Hecke (2021), buckled beams Sirote-Katz et al. (2024); Liu et al. (2024), and biholar sheets Ding and van Hecke (2022); El Elmi and Pasini (2024). Yet, even in this realm where one has complete control over the material architecture, no system has been able to realize the full generality of Eq. 1.
Here, we translate this abstract model into a physical design platform. We show how mechanical hysterons built from rigid bars and linear springs can realize the full generality of Eq. 1 by way of strong, tunable interactions that are ferromagnetic-like or anti-ferromagnetic-like. We then target three specific computations on sequences of mechanical inputs, and build the corresponding systems. Our results provide a replicable, general platform for processing mechanical information in materia.
Hysteron design. Figure 1A illustrates the basic bistable unit in our system: a rigid bar that is free to rotate about a central pivot. The bar is corralled between two posts that restrict its angle to an interval, $\theta^{-},\theta^{+}]$, with $\theta=0$ corresponding to vertical. A spring of stiffness $k_{1}$ and rest length $x_{0}$ attaches to the bar a distance $L$ from the pivot; the other end of this spring attaches to a long steel rod that translates freely along its axis via two linear bearings. The $x$-position of this rod, $\gamma$, serves as the global mechanical drive. All springs used in the experiments behave linearly, which we verify by measuring their force-versus-displacement curves when pulling up on a known mass resting on an electronic balance (Fig. 1B). We analyze these data to measure $k$ and $x_{0}$ for each spring used in the experiment.
We first consider the stability of a single rotor. Starting with the driving rod on the far left and gradually moving it to the right, the rotor begins in the (-) state where $\theta=\theta^{-}$, but becomes unstable and jumps to the (++) state ($\theta=\theta^{+}$) when the line running through the spring passes the pivot point of the rod. This transition occurs when the horizontal displacement of the driving rod, $\gamma$, exceeds a threshold $\gamma^{+}=-y\tan\theta^{-}$. Likewise, the rotor flips to ()(-) when $\gamma<\gamma^{-}=-y\tan\theta^{+}$. The rotor is thus a hysteron: it is monostable outside the interval $\gamma^{-},\gamma^{+}]$ and it is bistable within (Fig. 1C). We may move this interval of stability by shifting the position $x_{i}$ where hysteron $i$ couples to the driving rod (where $x_{i}=0$ corresponds to no shift). Supplementary Movie 1 shows these basic behaviors for a pair of uncoupled rotors.
Next, we create interactions by coupling rotors $i$ and $j$ together with additional springs of stiffness $k_{ij}$ and rest length $x_{0}$, mounted at a distance $\ell$ from the pivot point (Fig. 1A). Uncrossed coupling springs create cooperative (ferromagnetic-like) interactions, which encourage two hysterons to occupy the same state, forming the composite states -- or ++++ more readily. Crossed springs create frustrated (antiferromagnetic-like) interactions, which encourage them to occupy the states ++- or +-+ (Fig. 1D).
These interactions may be understood by torque balance. Suppose two rotors are held fixed at $\theta_{1}=\theta_{2}=0^{\circ}$. Coupling the rotors with two identical springs leads to zero net torque on rotor $1$, as the torques from the two springs cancel whether they are crossed or uncrossed. Moving rotor $2$ into the ++ state increases the distance spanned by one coupling spring and decreases the distance spanned by the other. Since these springs attach to rotor $1$ on different sides of the pivot, this change in the torque adds up—two positive numbers for uncrossed springs (pushing rotor 1 towards the ++ state) and two negative values for crossed springs (pushing rotor 1 towards the - state). Through these interactions, the switching thresholds for one hysteron thus become dependent on the state of the other, giving rise to an expanded set of thresholds. The ensuing complexity and rapid tunability is demonstrated in Supplementary Movie 2, where we demonstrate a multitude of transitions for two cooperatively-coupled hysterons.
Tunable interactions. The position where the coupling springs attach, $\ell$, offers a convenient way to tune the interaction between the hysterons continuously. To demonstrate this control, we perform experiments on two hysterons with a cooperative interaction, where we fix all other system parameters and vary $\ell$. We measure the switching thresholds by holding one rotor fixed in the $(+)$ or ()(-) state, and quasistatically moving the driving rod until the other hysteron undergoes a transition. For example, $\gamma^{+}_{1}(s_{2})$ denotes the measured threshold for hysteron 1 to switch to $(+)$ when hysteron 2 is in state $s_{2}$. Figure 2A shows the measured values of the eight transitions for this system, at four values of $\ell$.
When $\ell=0$, the coupling springs have no lever arm to exert a torque on either rotor  [^1] In practice we remove the coupling springs for the case $\ell=0$.. Here there are only four distinct thresholds, due to the degeneracies $\gamma_{i}^{\pm}(+)=\gamma_{i}^{\pm}(-)$. The particular order of the thresholds, $\gamma^{-}_{1}<\gamma^{-}_{2}<\gamma^{+}_{2}<\gamma^{+}_{1}$, gives rise to a set of stable states and transitions that we may represent in a transition graph Mungan et al. (2019); Bense and van Hecke (2021) – a directed graph that shows the transitions out of each stable state under increasing or decreasing drive (Fig. 2B). For $\ell=1.6$ cm, the thresholds split into eight distinct values. Nevertheless, the same stable states and transitions remain at $\ell=1.6$ cm and at $\ell=3.2$ cm.
Increasing $\ell$ to $4.6$ cm, two pairs of thresholds switch their ordering: $\gamma_{2}^{+}(+)$ falls below $\gamma_{2}^{-}(-)$, and $\gamma_{1}^{+}(+)$ falls below $\gamma_{2}^{+}(-)$. The first switch does not change the transition graph for the system, but the second does. Starting in --, the first transition upon increasing $\gamma$ is to +-+ at $\gamma=\gamma_{2}^{+}(-)$. But this state is now unstable, as $\gamma_{2}^{+}(-)>\gamma_{1}^{+}(+)$, meaning that hysteron 1 is above threshold to flip to $(+)$ as soon as hysteron 2 enters the $(+)$ state. This “vertical avalanche” van Hecke (2021) is denoted by the arrow from -- to ++++ in the second transition graph in Fig. 2B.
This tunability of the thresholds and interactions suggests that our mechanical hysterons should be able to exhibit a wide variety of behaviors. We confirm this notion in Supplementary Movies 2-5, where we demonstrate all possible transition graphs for two hysterons with cooperative interactions ($J_{ij}>0$) and frustrated interactions ($J_{ij}<0$) van Hecke (2021). This suite of behaviors includes a 2-bit analog-to-digital converter (Fig. S1A, transition graph $ix$). We provide further descriptions of these experiments in the Supplementary Information and in Fig. S1.
Kinematic model. We can use torque balance to connect the system parameters to the switching thresholds observed in the experiment. First, we assume the system is in a stable state $S$ that specifies the $\{s_{j}\}$. It is then straightforward to write down the torque $\tau_{ij}(s_{i},s_{j})$ of rotor $j$ on rotor $i$, in terms of the properties and mounting positions of the coupling springs. Similarly, one can compute the torque $\tau_{i}(\gamma,s_{i})$ on rotor $i$ by its driving spring. Exact expressions for $\tau_{i}(\gamma,s_{i})$ and $\tau_{ij}(s_{i},s_{j})$, which take into account the finite rest length of the springs, are provided in the Supplementary Information. Hysteron $i$ will undergo a transition when the sum of the torques on it vanish  [^2] To ensure this is an unstable equilibrium, it is necessary that $y>0$, but also that the driving spring $k_{i}$ is sufficiently strong compared to any coupling springs that might disrupt this bistability.:
$0=\tau_{i}(\gamma,s_{i})+\sum_{i\neq j}\tau_{ij}(s_{i},s_{j})\ .$ (2)
The $\gamma$ that solve this equation are the switching thresholds, $\gamma_{i}^{\pm}(S)$, wherein hysteron $i$ switches out of the state $s_{i}=\mp 1$  [^3] We do not presently have a precise condition on the parameters that would prohibit additional stable equilibria within the interval $(\theta_{i}^{-},\theta_{i}^{+})$.. We use Eq. 2 to compute the predicted thresholds for Fig. 2A and find good agreement with the experiments. Moreover, the model allows us to anticipate precisely where the qualitative behaviors of the system will change, e.g., the change in the transition graph at $\ell=4.18$ (Fig. 2A,B)  [^4] Another crossing of switching thresholds occurs at $\ell=4.90$; above this $\ell$ the avalanche from -- to ++++ occurs via the intermediate state ++-, as now $\gamma_{1}^{+}(-)<\gamma_{2}^{+}(-)$..
Linear interactions at arbitrary rotor angles. Equation 1 posits that the hysteron interactions are pairwise and linear. We now show that our system exhibits such interactions when the rest length $x_{0}$ of each driving spring is much less than $L+y$. In this limit, the torque on rotor $i$ by its driving spring is: $\tau_{i}=Lk_{i}\cos\theta_{i}(\gamma+x_{i}+y\tan\theta_{i})$. We emphasize that this result is for arbitrary $\theta_{i}$ and $\gamma$ as detailed in the Supplementary Information. Plugging this expression into Eq. 2 and solving for $\gamma$ gives the transitions $\gamma^{\pm}_{i}(S)$ for rotor $i$ when the system is in microstate $S$. The interaction between hysterons $i$ and $j$ may then be calculated by the difference, $\gamma_{i}^{\pm}(S|_{s_{j}=-})-\gamma_{i}^{\pm}(S|_{s_{j}=+})=2J_{ij}$, which yields:
$J_{ij}^{\pm}=\frac{1}{2Lk_{i}\cos\theta_{i}^{\mp}}(\tau_{ij}(\mp,+)-\tau_{ij}(\mp,-))\ ,$ (3)
independent of the state of all other rotors. Such a result establishes that the interactions are pairwise and linear, even for arbitrary rotor angles and finite coupling spring rest lengths. This linearity may be attributed to two key features of our system: (i) the affine relationship between the drive $\gamma$ and its torque on a rotor, and (ii) the additivity of the torques from the coupling springs.
Equation 3 additionally says that there are two sets of interactions strengths: $J_{ij}^{+}$ describes the interactions for hysteron $i$ when $\theta_{i}=\theta_{i}^{-}$, whereas $J_{ij}^{-}$ is the relevant matrix when $\theta_{i}=\theta_{i}^{+}$. Such abstract hysteron models have been proposed recently Szulc et al. (2022), where $J_{ij}$ in Eq. 1 is replaced by $J_{ij}^{\pm}$. In Fig. S2 we show that the simpler scenario, $J_{ij}^{+}=J_{ij}^{-}=J_{ij}$, occurs when the stopping posts are positioned symmetrically for each rotor, meaning $\theta_{i}^{+}=-\theta_{i}^{-}$ for each $i$.
We can also calculate the bare thresholds, $\gamma_{i,0}^{\pm}$, in terms of the torques on rotor $i$. They are found from the relation, $\gamma^{\pm}_{i}(S^{+})+\gamma^{\pm}_{i}(S^{-})=2\gamma_{i,0}^{\pm}$, where $S^{\pm}$ is the state with $s_{j}=\pm 1$ for all $j\neq i$. Using Eq. 2 and the expression for $\tau_{i}$ above yields:
$\gamma_{i,0}^{\pm}=-x_{i}-y\tan\theta_{i}^{\mp}-\frac{1}{2Lk_{i}\cos\theta_{i}^{\mp}}\sum_{j\neq i}(\tau_{ij}(\mp,+)+\tau_{ij}(\mp,-))\ .$ (4)
We emphasize that the sum in Eq. 4 represents a shift of the bare threshold due to the presence of interactions. Physically, it arises because the torque exerted by rotor $j$ on rotor $i$, averaged over the two states $s_{j}=\pm 1$, may be non-zero. By contrast, the interactions $J_{ij}$ measure the positive and negative departures from this mean, which are then added on top of this bare threshold, following Eq. 1.
Mapping to hysteron model at small rotor angles. We can gain a deeper understanding of the interactions by studying a limit where the rotor angles are small, $|\theta_{i}^{\pm}|\ll 1$. To emphasize the core relationships in this mapping, we specialize to the case where the two springs connecting rotors $i$ and $j$ have identical spring constants $k_{ij}$ and rest lengths $x_{0}$. To simplify expressions, we consider coupling springs with rest lengths that are small compared to either the rotor spacing or the mounting distance on the rotor: $x_{0}\ll w$ or $x_{0}\ll\ell$. We show in the Supplementary Information that under these conditions, the torque of hysteron $j$ on $i$ has a simple form: $\tau_{ij}=2g_{ij}k_{ij}\ell^{2}(\theta_{j}-\theta_{i})$, where the geometric factor $g_{ij}$ is defined by:
$g_{ij}=\begin{cases}1&\text{parallel springs,}\\ -1&\text{crossed springs.}\end{cases}$ (5)
Plugging this torque into Eqs. 3 and 4 at small angles, we obtain:
$J_{ij}$ $=$ $g_{ij}\frac{k_{ij}}{k_{i}}\frac{\ell^{2}}{L}(\theta_{j}^{+}-\theta_{j}^{-})\ ,$ (6)
$\gamma_{i,0}^{\pm}$ $=$ $-x_{i}-y\tan\theta_{i}^{\mp}-\sum_{j\neq i}\frac{k_{ij}}{k_{i}}\frac{\ell^{2}}{L}g_{ij}(\theta_{j}^{+}+\theta_{j}^{-}-2\theta_{i}^{\mp})\ ,$ (7)
where we retained $y\tan\theta_{i}^{\mp}$ in place of $y\theta_{i}^{\mp}$ in Eq. 7 so that it gives the exact threshold in the absence of coupling.
Figure S3 shows that the small-$\theta$ model of Eqs. 1, 6, and 7 provides a good approximation to the exact kinematic model, Eq. 2, for both cooperative and frustrated interactions. Moreover, Eqs. 6 and 7 provide important physical insights into our system. For instance, they tell us that only one interaction matrix is sufficient to describe the system at small rotor angles ($J_{ij}^{+}=J_{ij}^{-}=J_{ij}$), even without the additional symmetry of the rotor angles assumed in Fig. S2. They also identify an inherent non-reciprocity of the interactions that we explore in the next section. More valuable still, Eqs. 6 and 7 are a blueprint for designing the thresholds and interactions in our system at will—an enticing topic that comprises the remainder of this article.
Non-reciprocal interactions. Returning first to Fig. 2A, a closer study reveals that those interactions are non-reciprocal: The splitting of the blue curves, $\gamma_{2}^{\pm}(-)-\gamma_{2}^{\pm}(+)$, is larger than the splitting of the black curves, $\gamma_{1}^{\pm}(-)-\gamma_{1}^{\pm}(+)$. The non-reciprocity of the interaction can immediately be seen via Eq. 6, where $J_{ij}$ depends on the angular interval $\theta_{j}^{+}-\theta_{j}^{-}$ but does not depend on $\theta_{i}^{\pm}$. The larger splitting of hysteron 2 in Fig. 2A is thus due to the larger angular interval of hysteron 1. To test this result quantitatively, Fig. 2C compares Eq. 6 with the value of $\gamma_{i}^{\pm}(-)-\gamma_{i}^{\pm}(+)$ from the full kinematic model and the experiments. The comparison shows that Eq. 6 captures the observed non-reciprocity, as well as the predicted $\ell^{2}$ scaling.
Having established a physical picture for the switching thresholds and interactions for pairs of mechanical hysterons, the remainder of this article is devoted to targeting more complex behaviors with two, four, and many coupled hysterons.
Hysteron latching. Holding a blueprint for frustrated non-reciprocal interactions allows one to target an exotic behavior termed “latching”. This behavior was proposed by Lindeman et al. Lindeman et al. (2025) to explain how amorphous solids form multiple memories of asymmetric driving. Here we demonstrate it with our mechanical hysterons in Fig. 3A and Supplementary Movie 6.
Starting in the -- state at $\gamma=0$, we drive the system to $\gamma=10$ cm, which puts the hysterons in the ++- state. This state remains stable as we return $\gamma$ back to 0; hysteron 1 is now “latched” in the $(+)$ state. A larger driving amplitude to $\gamma=32$ cm and back to 0 releases hysteron 1 back to ()(-) via a sequence of intermediate states: $-\rightarrow++\rightarrow-+$$\rightarrow--$.
Latching necessarily violates return-point memory Barker et al. (1983); Sethna et al. (1993); Keim et al. (2019), which would stipulate that the same state would recur each time the driving returns to its common endpoint, $\gamma=0$. Indeed, Lindeman et al. Lindeman et al. (2025) showed that latching requires a strong, frustrated, non-reciprocal interaction: $J_{12}<J_{21}<0$. Figure 3B shows that our setup satisfies this requirement. We obtain this frustrated non-reciprocal interaction by confining hysteron 1 to a smaller angular interval than hysteron 2 and coupling them with crossed springs. The full requirements for latching, and how we target them using our mechanical hysterons, is provided in the Supplementary Information.
Counting driving cycles. Frustrated interactions enable other computational behaviors, such as counting the number of mechanical driving cycles that have been applied. Our design uses an antiferromagnetic chain of hysterons with uniform frustrated interactions and uniform hysteresis, $\gamma^{+}_{i}-\gamma^{-}_{i}$. We alternately raise and lower both switching thresholds, $\gamma_{i}^{+}$ and $\gamma_{i}^{-}$, for each hysteron (Fig. 4A). This staggering breaks the degeneracy of the two ground states so that ++-++-... has a higher energy than +-++-+.... Consider now a stable state with one “domain wall” in the bulk where the two ground states meet. Both hysterons at the domain wall experience two interactions that cancel, lowering their barrier to change state via the global drive. This grows the ++-++-... domain by one site for each half-cycle of driving. The location of the domain wall thus encodes the number of driving cycles; $2n$ hysterons can record this number up to $n$.
This design is easily translated to our mechanical system by forming a chain of rotors that are coupled to their nearest neighbors with crossed springs, and attaching them to the driving rod with $x_{i}$ that alternate between negative and positive values (Fig. 4B). In particular, we set $x_{2m-1}=-L\sin(\theta_{0})$, $x_{2m}=L\sin(\theta_{0})$, where $\theta_{0}=\theta_{i}^{+}=-\theta_{i}^{-}$, although other values can also work (see the Supplementary Information). Finally, we set $x_{1}=-3L\sin(\theta_{0})$ at the start of the chain to nucleate a domain wall there. Supplementary Movie 7 and Fig. 4B show the system evolution, starting from the higher-energy ground state. Once the domain wall reaches the end of the chain, the configuration is stable to further driving.
Counting modulo 2. Figure 5A shows a system of four hysterons that, when driven at unit amplitude, produces a steady state that takes two cycles to traverse. It concludes even cycles in the state --+-+, and it concludes odd cycles in the state --++- (Fig. 5B). The system thus performs a basic computation: discriminating between even and odd numbers of driving cycles, i.e., counting modulo 2. Such a multiperiodic response has been identified in simulated hysteron systems Keim and Paulsen (2021); Lindeman and Nagel (2021); Szulc et al. (2022). Here we realize it in the lab.
We start by translating the sign and relative strength of the interactions in Fig. 5A into a choice of springs and their mounting points on the rotors. One pair of springs spans a longer distance from rotor 1 to rotor 3; we attach a length of steel wire to extend these springs while staying within their linear response. We then set $\theta_{i}^{+}-\theta_{i}^{-}$ for each hysteron to reflect its hysteresis in the design. To ensure constant-amplitude driving, we use the linear bearings that hold the driving rod as a “sandwich” that sets the endpoints of the driving.
The mounting positions $x_{i}$, the angular intervals $\theta_{i}^{+}-\theta_{i}^{-}$, the coupling positions $\ell_{ij}$, and the two endpoints of the driving offer a total of 14 continuous degrees of freedom. We also allow ourselves to mount some of the springs within a pair at unequal $\ell_{ij}$. To target the sequence of states in Fig. 5B, we drive the system to observe its behavior and then adjust one or more of these degrees of freedom to try to obtain more of the desired sequence. This iterative process eventually produced the configuration shown in Fig. 5C; its behavior under cyclic driving is shown in Supplementary Movie 8. To show the resulting steady state, we measure the value of $\gamma$ where each transition occurs, and we plot the sequence of states as a function of time in Fig. 5D. The sequence matches that of Fig. 5B, it takes two periods to traverse, and it repeats indefinitely, as desired  [^5] There is one subtle difference in the pathways: The experiment navigates the downward avalanche via the unstable intermediate state +-+++-, whereas the design in Fig. 5A passes through ++-++- instead. We found that we could match the experiment by changing the bare threshold $\gamma_{1,0}^{-}$ for the top hysteron in Fig. 5A from 0 to $0.85$..
Discussion. Recent work has demonstrated other mechanical designs that can count modulo two ten Wolde and Farhadi (2024), count driving cycles up to $n$ Kwakernaak and van Hecke (2023), or encode an analog mechanical input into a binary string Hyatt and Harne (2023). Although those systems use bistable mechanical elements, each one uses a different strategy to navigate the pathways between its states. In contrast, we have presented a general platform that can be reconfigured to perform all of the above functions, plus a latching behavior that was recently identified in hysteron simulations Lindeman et al. (2025). This platform could find use in the design of mechanisms that navigate through precise sequences of states, e.g., in robotics Kamp et al. (2024). More broadly, our work identifies a vast potential for designed materials that inherently engage with memory Keim et al. (2019), as their current state is a by-product of their driving history along with their overall design, which our platform now brings under the experimenter’s control.
One principle of our approach was to use nonlinearities sparingly, so that the behaviors of the system could be rationalized as simply as possible. Nevertheless some important questions remain: Schrecengost Schrecengost (2025) showed that mixed interactions ($J_{ij}J_{ji}<0$) can be ruled out in a certain region of parameter space, but a full treatment of this question requires further work. Other designs for interacting mechanical hysterons are also being proposed; Shohat and van Hecke Shohat and van Hecke (2025) use theory and simulations to explore two-dimensional networks of bilinear hysterons as a route to exotic pathways, including mixed interactions and multiperiodic cycles. Understanding what are the general constraints on $\gamma_{i,0}^{\pm}$ and $J_{ij}$ for any realizable structure or metamaterial is an important open question Baconnier et al. (2025).
Designing the response of larger collections of mechanical hysterons presents another challenge Teunisse and van Hecke (2025); Muhaxheri et al. (2025). In Fig. 5 we obtained a period-2 response by aiming for a set of target hysteron parameters that were originally found via a random search Keim and Paulsen (2021). Our approach for dialing in that precise set of couplings was to iteratively adjust the mounting positions of springs and the post angles, to pick up more and more of the sought after multiperiodic cycle. This process is reminiscent of supervised learning Stern and Murugan (2023); Altman et al. (2024), suggesting that learning might be a viable approach for targeting transition graphs at large. Our system offers a wealth of possible learning degrees of freedom, as there are numerous geometric parameters that can be tuned continuously to affect the switching thresholds, including $x_{i}$, $\theta_{i}^{\pm}$, $\ell_{ij}$, and $L_{i}$. Further work is needed to determine whether a suitable learning algorithm could be used to navigate between different transition graphs, or classes of transition graphs with a desired functionality.
Methods. Rotor hysteron construction. We construct each rotor hysterons from two 1/21/2”-diameter optical posts attached in a T configuration, which we mount in a radial steel ball bearing affixed to an optical table. We remove the bearing grease to further lower their friction to rotation. We use steel extension springs to couple the rotors to one another and to a 1/2”-diameter steel shaft that is mounted in two linear bearings. This 72”-long rod translates freely along the horizontal direction and provides the global mechanical drive, $\gamma$. We define $\gamma$ as the horizontal displacement of this rod in centimeters, which we measure with a precision of $0.1$ cm via a ruler affixed under the rod. We define $x_{i}$, the mounting position of driving spring $i$ on the driving rod, such that $x_{i}=0$ corresponds to no shift when the driving rod is at $\gamma=0$.
Spring stiffnesses. We measure the spring stiffnesses from force-versus-displacement data that we obtain by anchoring the springs to a known mass on an electronic balance, and recording the force on the balance as the top of the spring is extended and then retracted (Fig. 1B). These measurements show that the springs behave linearly over the range of extensions used in the experiments.
Abstract hysteron simulations. We verified the period-2 behavior of Fig. 5A using the open-source hysteron software package Keim and Paulsen (2020).
Acknowledgements. I thank Nathan Keim for suggesting how to apply a global drive, for providing the parameters used in Fig. 5A, and for many fruitful discussions. I am grateful to Zachariah Schrecengost for identifying how the interactions could be treated at arbitrary rotor angles. I thank Vidyesh Anisetti, Eadin Block, Christian Santangelo, and Jennifer Schwarz for contributing to early hysteron designs, and I thank Martin van Hecke, Nathan Keim, Chloe Lindeman, Gentian Muhaxheri, and Dor Shohat for comments on the manuscript. This work was supported by Syracuse University research subsidy funds and was initiated at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452.
Supplementary Information for
“Mechanical hysterons with tunable interactions of general sign”
Joseph D. Paulsen
I Targeting specific transition graphs
We start by demonstrating four distinct transition graphs in a sequence where we alter only the mounting position of driving spring 2, thereby changing only $x_{2}$. Supplementary Movie 2 shows this sequence of behaviors, which correspond to transition graphs $i$-iviv in Fig. S1A. The experimental parameters are tabulated in Fig. S1C.
To understand how the system moves between these four transition graphs, we plug these parameters into the kinematic model, calculate the switching thresholds, and plot them in Fig. S1B. The switching thresholds of hysteron 1 are constant throughout this sequence, as they depend on the state of hysteron 2 but not its driving. The thresholds for hysteron 2 move in concert without changing their spacing, as changing $x_{2}$ shifts each threshold by $x_{2}$. Careful inspection of the switching thresholds in Fig. S1B shows that they indeed produce graphs $i$-iviv in Fig. S1A, faithful to the experiment.
Next, we access transition graph $v$ by decreasing $\Delta\theta_{1}$, increasing $\Delta\theta_{2}$, and changing $x_{2}$ (Supplementary Movie 3). We access transition graph $vi$ by setting $\Delta\theta_{1}$ = $\Delta\theta_{2}$ and changing $x_{2}$ once again (Supplementary Movie 4). This exhausts all possible transition graphs for two hysterons with cooperative interactions van Hecke (2021).
Three additional transition graphs can be obtained by using frustrated interactions van Hecke (2021); they are graphs viivii-ixix in Fig. S1A. Supplementary Movie 5 shows how we can realize these graphs in the experiment while varying only $x_{2}$ to navigate between them. All three graphs feature at least one “horizontal avalanches”. To understand how they occur, start in the -- state in graph viivii. The corresponding thresholds in Fig. S1B show that upon increasing $\gamma$, hysteron 2 switches first, as $\gamma_{2}^{+}(-)<\gamma_{1}^{+}(-)$. Now in the +-+ state, the next event is at $\gamma=\gamma_{1}^{+}(+)$ into ++++. But this state is not stable, as this $\gamma$ is below $\gamma_{2}^{-}(+)$, which brings the system immediately to ++-. An analogous process underlies each of the $\leftrightarrow+-$ avalanches in graphs viivii-ixix.
Among these graphs, perhaps the most exotic is $ix$, which navigates the sequence: --, +-+, ++-, ++++ under unidirectional driving. This behavior constitutes a 2-bit analog-to-digital converter van Hecke (2021). We now describe the process used to target this transition graph, to illustrate how behaviors may be rationally designed in our system. First, we need a strong coupling between the global drive and hysteron 1, so that it is ()(-) at small $\gamma$ and $(+)$ at large $\gamma$. We also need a strong frustrated interaction between hysterons 1 and 2 to navigate the $\leftrightarrow+-$ avalanches. In particular, the interaction of hysteron 1 on hysteron 2 must open up a central interval, $\gamma_{2}^{+}(-)<\gamma<\gamma_{2}^{-}(+)$ where the state of hysteron 2 depends only on the current state of hysteron 1. Figure S1D illustrates this condition, and an examination of graphs viivii-ixix in Fig. S1 shows that they satisfy it. We target this condition by selecting a large $k_{12}$ and a sufficiently large $\theta_{1}^{\pm}$; we find that the same $\pm 38.8^{\circ}$ used in the other designs works here too. Finally, a weak coupling of hysteron 2 to the global drive enables a transition into the ++++ state at large $\gamma$ and into -- at small $\gamma$. These considerations lead to a tentative design for the spring constants: $k_{1}\approx k_{12}\gg k_{2}$. Here we use $k_{1}=k_{12}=0.46$ N/cm and $k_{2}=0.07$ N/cm, which we found to suffice.
Two additional transition graphs are possible for “mixed” interactions, where the signs of $J_{ij}$ and $J_{ji}$ differ van Hecke (2021). We do not see a way to realize such interactions in our system.
II Torque from the driving rod
Here we derive the torque exerted by the driving rod on a hysteron. A driving spring of stiffness $k_{i}$ and rest length $x_{0}$ mounts to rotor $i$ at a position $L>0$ from the fulcrum of the rotor, with coordinates $\vec{r}_{L}=\{L\sin\theta_{i},L\cos\theta_{i}\}$. The other end of the driving spring connects to the driving rod at coordinates $\vec{r}_{\text{bar}}=\{\gamma+x_{i},-y\}$. The springs in the experiment obey Hooke’s Law: $F=k_{i}(x-x_{0})$, where $x=|\vec{r}_{\text{bar}}-\vec{r}_{L}|$. In vector form,
$\vec{F}=k(\vec{r}_{\text{bar}}-\vec{r}_{L})\left(1-\frac{x_{0}}{|\vec{r}_{\text{bar}}-\vec{r}_{L}|}\right)\ .$ (S1)
The torque on rotor $i$ is $\tau_{i}=-\vec{r}_{L}\times\vec{F}$, where use a minus sign when defining the torque, so that clockwise rotation of the rotors corresponds to increasing $\theta$. Computing the cross product by components, we obtain the exact expression:
$\tau_{i}=Lk_{i}\cos\theta_{i}(\gamma+x_{i}+y\tan\theta_{i})\left(1-\frac{x_{0}}{\sqrt{(\gamma+x_{i}-L\sin\theta_{i})^{2}+(y+L\cos\theta_{i})^{2}}}\right)\ .$ (S2)
We use Eq. S2 when computing switching thresholds in the kinematic model.
When the rest length of the driving spring, $x_{0}$, is small compared to $L+y$, the torque is approximately:
$\tau_{i}=Lk_{i}\cos\theta_{i}(\gamma+x_{i}+y\tan\theta_{i})\ .$ (S3)
We use Eq. S3 to derive Eqs. 3 and 4 in the main text and for the small-$\theta$ model.
III Torques between coupled hysterons
Here we derive expressions for the torque exerted by one hysteron on another. We begin with the most general case: arbitrary angles for the rotors, finite rest lengths for the springs, and arbitrary mounting positions of the springs on the rotors (Eq. S5). We then expand this result to leading order in $\theta_{1},\theta_{2}$ (Eq. S6). We then specialize this small-angle result to the “paired springs” that are used exclusively in the main text. The result for the “parallel” configuration turns out to be independent of the rest length of the springs (Eq. S7), whereas the “crossed” configuration requires one more limit to reach a similarly simple expression (Eq. S9).
III.1 Single coupling spring
We consider a single coupling spring of stiffness $k$ and rest length $x_{0}$ that mounts to rotors $1$ and $2$ at positions $\ell_{1}$ and $\ell_{2}$ from the fulcrums of the rotors, respectively, where the signs of $\ell_{1}$, $\ell_{2}$ indicate the side of the fulcrum where they attach. The coordinates of the mounting position on rotor $1$ are $\vec{r}_{1}=\{\ell_{1}\sin\theta_{1},\ell_{1}\cos\theta_{1}\}$, and on rotor $2$ they are $\vec{r}_{2}=\{w+\ell_{2}\sin\theta_{2},\ell_{2}\cos\theta_{2}\}$. The springs in the experiment obey Hooke’s Law: $F=k(x-x_{0})$, where $x=|\vec{r}_{2}-\vec{r}_{1}|$. In vector form,
$\vec{F}=k(\vec{r}_{2}-\vec{r}_{1})\left(1-\frac{x_{0}}{|\vec{r}_{2}-\vec{r}_{1}|}\right)\ .$ (S4)
The torque on rotor $1$ is $\tau=-\vec{r}_{1}\times\vec{F}$, where use a minus sign when defining the torque, so that clockwise rotation of the rotors corresponds to increasing $\theta$. Computing the cross product by components, we obtain the exact expression:
$\tau=k\ell_{1}(w\cos\theta_{1}+\ell_{2}\sin(\theta_{2}-\theta_{1}))\left(1-\frac{x_{0}}{\sqrt{(w+\ell_{2}\sin\theta_{2}-\ell_{1}\sin\theta_{1})^{2}+(\ell_{2}\cos\theta_{2}-\ell_{1}\cos\theta_{1})^{2}}}\right)\ .$ (S5)
Expanding Eq. S5 in $\theta_{1}$ and $\theta_{2}$, we obtain the torque on hysteron 1 to leading order in $\theta_{1},\theta_{2}$:
$\tau=k\ell_{1}(w+\ell_{2}(\theta_{2}-\theta_{1}))\left(1-\frac{x_{0}}{L_{0}}\right)+k\ell_{1}(\ell_{2}\theta_{2}-\ell_{1}\theta_{1})\frac{w^{2}x_{0}}{L_{0}^{3}}+$...\ , (S6)
where $L_{0}=\sqrt{(\ell_{2}-\ell_{1})^{2}+w^{2}}$ is the distance spanned by the spring when $\theta_{1}=\theta_{2}=0$.
Equations S5 and S6 apply when rotor $2$ is to the right of rotor $1$. If rotor $2$ is on the left, sending $\ell_{1}\rightarrow-\ell_{1}$ and $\ell_{2}\rightarrow-\ell_{2}$ gives the relevant formulas, as this represents a $180^{\circ}$ rotation about the center of rotor $1$, which leaves both rotor angles and the coupling spring length unchanged.
III.2 Paired springs at small rotor angles
Using Eq. S6, it is straightforward to write down the net torque $\tau_{12}=\tau+\tau^{\prime}$ on rotor 1 for the general case where a second spring of stiffness $k^{\prime}$ and rest length $x_{0}^{\prime}$ is mounted at distinct positions $\ell_{1}^{\prime}$ and $\ell_{2}^{\prime}$. Pairing the springs as in Fig. 1D with spring constants $k=k^{\prime}$ and $x_{0}=x_{0}^{\prime}$ greatly simplifies the form of the interaction. Mounting the springs in a “parallel” configuration, $\ell=\ell_{1}=\ell_{2}=-\ell_{1}^{\prime}=-\ell_{2}^{\prime}$, we find:
$\tau_{12}=2k\ell^{2}(\theta_{2}-\theta_{1})\ .$ (S7)
Notably, the net torque does not depend on the finite rest length of the springs, $x_{0}=x_{0}^{\prime}$, or the distance between rotors, $w$, due to the symmetry of such paired coupling springs.
Mounting instead in a “crossed” configuration, $\ell=\ell_{1}=-\ell_{2}=-\ell_{1}^{\prime}=\ell_{2}^{\prime}$, we obtain:
$\tau_{12}=-2k\ell^{2}(\theta_{2}-\theta_{1})-2k\ell^{2}\frac{x_{0}}{\sqrt{4\ell^{2}+w^{2}}}\left(2\theta_{1}-\frac{4\ell^{2}}{4\ell^{2}+w^{2}}(\theta_{2}+\theta_{1})\right)\ .$ (S8)
Specializing to either $x_{0}\ll w$ or $x_{0}\ll\ell$ simplifies this expression to:
$\tau_{12}=-2k\ell^{2}(\theta_{2}-\theta_{1})\ ,$ (S9)
which is equal and opposite to Eq. S7.
IV Targeting a latching behavior
In Fig. 3 of the main text we demonstrate a “latching” behavior, which had been introduced in the context of sheared amorphous solids in recent work by Lindeman et al. Lindeman et al. (2025). Here we describe how to rationally target this behavior using our mechanical hysterons. Lindeman et al. Lindeman et al. (2025) showed how latching is tied to the ordering of the switching thresholds. They found that latching requires:
$\gamma_{1}^{-}(-)<0<\gamma_{2}^{-}(-)<\gamma_{2}^{-}(+)<\gamma_{1}^{-}(+)\ ,$ (S10a)
$\gamma_{1}^{+}(-)<\gamma_{2}^{+}(-)\ .$ (S10b)
We now show how we translate this ordering into a concrete design in our mechanical system. We start by noting that Eq. S10a implies: $0<\gamma_{2}^{-}(+)-\gamma_{2}^{-}(-)<\gamma_{1}^{-}(+)-\gamma_{1}^{-}(-)$, which we rewrite as: $0<-J_{21}<-J_{12}$. We obtain this non-reciprocal frustrated interaction by confining hysteron 1 to a smaller angular interval than hysteron 2 and coupling them together with crossed springs following Eq. 6. Next, we note that adjusting $x_{2}$ moves all thresholds for hysteron 2 in concert without altering their spacing. We thus move $x_{2}$ until $\gamma_{2}^{-}(+)<\gamma_{1}^{-}(+)$ and $\gamma_{1}^{-}(-)<\gamma_{2}^{-}(-)$. Then, we fulfill $\gamma_{1}^{-}(-)<0<\gamma_{2}^{-}(-)$ by moving $x_{1}$ and $x_{2}$ together by equal amounts until those thresholds straddle $\gamma=0$. Finally, $\gamma_{1}^{+}(-)<\gamma_{2}^{+}(-)$ turns out to be satisfied already through the larger angular interval of hysteron 2.
V Counting driving cycles
Here we provide additional details regarding the accumulator design in the main text (Fig. 4). The basic idea of our design can be verified on a computer Keim and Paulsen (2020) by applying cyclic driving of any fixed amplitude $\gamma_{0}$, setting thresholds $\gamma_{2m-1}^{+}>-\gamma_{2m-1}^{-}=\gamma_{0}/2$ for the odd hysterons and $\gamma^{\pm}_{2m}=-\gamma^{\mp}_{2m-1}$ for the even hysterons (where $m=1,2,3$...), and setting nearest-neighbor interactions of strength $J_{i,i+1}=-\gamma_{0}$. The thresholds for hysteron 1 should be raised by $\gamma_{0}$ to allow a domain wall to nucleate there.
More precisely, our accumulator design detects cycles of intermediate amplitude. Small-amplitude driving $\gamma(t)$ that is bounded by $\max(\gamma_{i}^{-})<\gamma(t)<\min(\gamma_{i}^{+})$ will not change the state of any hysteron, whereas large-amplitude driving that overpowers the frustrated interactions in the bulk can send the system to its final absorbing state. Intermediate-amplitude cycles advance the domain wall by two hysterons (one per half-cycle).
Figure S1: Suite of behaviors for an interacting hysteron pair. A. Transition graphs obtained in the experiment, showing all states and transitions accessible from the initial state, -- (Supplementary Movies 2-5). Graphs $i$ and iiii are “Preisach” graphs that do not require interactions van Hecke (2021), whereas graphs iiiiii-vivi require cooperative interactions. To show how one can move between different transition graphs, we demonstrate them all using cooperative interactions, without changing the driving or coupling springs (“Configuration 1” in C). Graphs viivii-ixix require frustrated interactions; we obtain them with a second set of driving and coupling springs (“Configuration 2” in C). B. Switching thresholds calculated from the model, using the experimental parameters. C. Experimental parameters. We fix $x_{1}=0$ cm, $y=14.5$ cm, $L=6.3$ cm, $\ell=4.7$ cm, and $w=15.2$ cm. D. Schematic illustrating the role of interactions on hysteron 2 in graphs $i$-ixix. Top row: hysteron 1 is $(+)$; bottom row: hysteron 1 is ()(-). Smaller cooperative interactions are required to create graphs $i$-vivi in B, whereas larger frustrated interactions are needed for graphs viivii-ixix.
Figure S2: Diagram showing equality of $J_{ij}^{+}$ and $J_{ij}^{-}$ when $\theta_{i}^{+}=-\theta_{i}^{-}$ and $\ell_{i}=-\ell_{i}^{\prime}$ for all $i$. A. Top schematic: Generic configuration with torque $\tau_{12}(+,-)$ on rotor $1$. Locations of both parallel and crossed springs are illustrated, but only one set is present at a time. If $\theta_{i}^{+}=-\theta_{i}^{-}$ and $\ell_{i}=-\ell_{i}^{\prime}$ for $i=1,2$, then reflecting this schematic about the horizontal is equivalent to flipping the states of both rotors, implying $\tau_{12}(-,+)=-\tau_{12}(+,-)$. B. Analogous pair of schematics showing $\tau_{12}(-,-)=-\tau_{12}(+,+)$ under the same conditions. Using Eq. 3 in the main text, we have $J_{12}^{+}=(\tau_{12}(-,+)-\tau_{12}(-,-))/C^{-}=(\tau_{12}(+,+)-\tau_{12}(+,-))/C^{+}=J_{12}^{-}$, where $C^{\mp}=2Lk_{i}\cos\theta_{i}^{\mp}$.
Figure S3: Cooperative and frustrated interactions. A. We use the same system parameters as in Fig. 2 and calculate switching thresholds from the kinematic model (solid lines, Eq. 2) and the small-$\theta$ model (dashed lines, Eqs. 1, 6, and 7). The small-$\theta$ model captures the gross behavior of the system in spite of the large $\theta_{i}$ involved. B. Same analysis but with crossed springs, producing a frustrated interaction ($J_{ij}<0$) so that now $\gamma_{i}^{\pm}(+)>\gamma_{i}^{\pm}(-)$.